Blue Angel, Inc., a private firm in the holiday gift industry, is considering a new project. The company currently has a target debt–equity ratio of .40, but the industry target debt–equity ratio is .35. The industry average beta is 1.40. The market risk premium is 7 percent, and the risk-free rate is 5 percent. Assume all companies in this industry can issue debt at the risk-free rate. The corporate tax rate is 34 percent. The project requires an initial outlay of $677,000 and is expected to result in a $97,000 cash inflow at the end of the first year. The project will be financed at the company’s target debt–equity ratio. Annual cash flows from the project will grow at a constant rate of 5 percent until the end of the fifth year and remain constant forever thereafter.

      

Calculate the NPV of the project.
Answer #1

Answer:-

Given
Debt / equity ratio = 0.40 =0.40 / 1
Debt / (equity + debt) = 0.40 / (0.40 + 1)
Debt / total capital = 0.4 / 1.4 = 0.29
Ie Total capital = equity +debt = 1.40
Equity / ( equity + debt ) = 1 / 1.4 = 0.71
Beta = 1.40
Risk free rate (RFR) = 5%
Market risk premium = 7%
Cost of equity = RFR + Beta ( Market risk premium)
Cost of equity = 5% + 1.4 ( 7%)
Cost of equity = 5% + 9.8% = 14.8 %
Cost of debt = Risk free rate = 5% (Given)
tax rate = 34%

WACC = cost of equity x wt of equity in total capital + cost of debt x wt. of debt in total capital x ( 1 - tax rate)
WACC = 14.8 % x 0.71 + 5% x 0.29 x ( 1 - 0.34)
WACC = 10.51% + 0.957 %
WACC = 11.47 % = 0.1147

The projects initial outlay ie. cash outflow = $ 677000

The cash inflows and their discounted values are below

Year 1 CF = $ 97000
Year 2 CF = $ 97000 x ( 1+0.5) = $ 97000 (1.05) = $101850
PV = $ 101850 / ( 1.1147) = $ 91369.87
Year 3 CF = $ 97000 x (1.05)2 = $ 106942.5
PV = $ 106942.5 / ( 1.1147)2 = $ 86066.54
Year 4 CF = $ 97000 x ( 1.05)3 = $ 112289.63
PV= $ 112289.63 / (1.1147)3 = $ 81865.2
Year 5 CF = $ 97000 x ( 1.05)4 = $ 117904.11
PV = $ 117904.11 / (1.1147)4 = $ 76365.45
Cash flows constant there after
The last term = $ 117904.11 / ( WACC - g) = $ 117904.11 / ( 0.1147 - 0.05) = $ 117904.11 / 0.647 = $ 182232

NPV = - $ 677000 + $ 97000 + $ 91369.88 + $ 86066.54 + $ 81865.2 + $ 76366.45 + $ 182232
NPV = $ 677000 + $ 614900
NPV = - $ 62100

Therefore the NPV of the project is - $ 62100 which is negative